• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name S. Sirca

    Log entry time 23:56:24 on April29,2008

    Entry number 235872

    keyword=Swing shift summary


    Swing shift summary | S. Sirca (SL), D. Armstrong (TO), C. Dutta (3rd)
    ======================================================================

    Events
    ------
    16:00 No beam upon start of shift (ongoing beam studies)
    17:45 Beam is back, target is Ta
    18:20 Insert C target
    18:50 Insert LH2 target
    18:55 BB MWDC HV trip (S16), recover in 10 minutes
    19:38 Start ramp down of HRSL momentum to 0.997 MeV/c for kinematics C-III
    19:51 Ready for production running in kinematics C-III
    20:30 Problems restoring HRSL HV (mainframe 7)
    21:10 Request hall access to reset HV crate
    22:10 Request beam again
    23:55 Beam away for PSS issues that can not be cleared

    Runs
    ----
    18:04 4387 3.0 uA on Ta (run to set prescales)
    18:08 4388 0.5M 3.0 uA on Ta
    18:23 4389 3.0 uA on C (run to set prescales, see HALOG 235822)
    18:28 4390 3.0 uA on C
    18:35 4391 0.5M 2.0 uA on C
    19:10 4392 3.5 uA on LH2 (run to set prescales)
    19:14 4393 3.5 uA on LH2 (run to set prescales)
    19:17 4394 3.5 uA on LH2 (run to set prescales)
    19:21 4395 3.5 uA on LH2
    19:52 4396 3.5 uA on LH2 (PS test)
    19:55 4397 3.5 uA on LH2
    20:09 4398 1.0M 3.5 uA on LH2
    20:24 4399 1.0M 3.5 uA on LH2
    20:39 4400 0.2M 3.5 uA on LH2 (NB start of HRSL HV problems)
    22:39 4401 0.7M 3.5 uA on LH2 (NB after HV problems resolved)
    22:43 4402 1.0M 3.5 uA on LH2
    22:58 4403 1.0M 3.5 uA on LH2
    23:10 4404 1.0M 3.5 uA on LH2
    23:25 4405 1.0M 3.5 uA on LH2
    23:39 4406 1.0M 3.5 uA on LH2 beam is off